# Unifying the Extremes: Developing a Unified Model for Detecting and Predicting Extremist Traits and Radicalization

### v1

@@ Abstract
The proliferation of ideological movements into extremist factions via social media has become a global concern. While radicalization has been studied extensively within the context of specific ideologies, our ability to accurately characterize extremism in more generalizable terms remains underdeveloped. In this paper, we propose a novel method for extracting and analyzing extremist discourse across a range of online community forums. By focusing on verbal behavioral signatures of extremist traits, we develop a framework for quantifying extremism at both user and community levels. Our research identifies 11 distinct factors, which we term “The Extremist Eleven,” as a generalized psychosocial model of extremism. Applying our method to various online communities, we demonstrate an ability to characterize ideologically diverse communities across the 11 extremist traits. We demonstrate the power of this method by analyzing user histories from members of the incel community. We find that our framework accurately predicts which users join the incel community up to 10 months before their actual entry with an AUC of $>0.6$, steadily increasing to AUC $\sim 0.9$ three to four months before the event. Further, we find that upon entry into an extremist forum, the users tend to maintain their level of extremism within the community, while still remaining distinguishable from the general online discourse. Our findings contribute to the study of extremism by introducing a more holistic, cross-ideological approach that transcends traditional, trait-specific models.
* equal contribution
@@ 1 Introduction
RQ1. Can an unsupervised method combining NLP techniques and psychosocial theories of extremism identify characteristics of extremism in online communities?
@@
RQ3. Can our framework reliably signal the likelihood of an individual’s eventual active engagement with an extremist community based on their evolving language patterns?
We aim to explore and understand how various psychological and social components contribute to — and characterize — radicalization across a range of extremist groups. 111 See §A.3 for semantic clarification. Given that numerous theories have been developed and applied to particular extremist groups and belief systems, we explore a method to unify them in a single model, allowing us to extract factors of extremism in broad, unsupervised data. To make use of existing psychometric and survey-based tools for measuring an individual’s tendency toward extremism at a large scale on social media data, we leverage natural language processing (NLP) techniques to obtain scores for the items using only the available text, requiring no manual labeling or supervision. We identify 11 orthogonal dimensions — the “Extremist Eleven” — that reflect important factors that can be used to broadly characterize key traits linked to extremist ideologies (RQ1). These dimensions are psychosocially meaningful and offer a more complete understanding of the cognitive and emotional drivers that underlie engagement with extremist groups. We explore the face validity of our method by examining how these factors emerge in communities known for extremist discourse and in a diverse representation of social media text that captures political discourse more broadly, as well as non-topic-specific posts.
Additionally, we demonstrate the utility of our methods and results in two preliminary ways. First, we show how the Extremist Eleven can be applied to characterize the psychological composition of both new and existing online communities, offering insights into the shared and unique traits that define these spaces (RQ2). Second, we find that we can accurately and reliably identify individuals who are likely to affiliate with and bind themselves to an extremist online community well in advance — more than 6 months prior to actively joining such communities (RQ3). This predictive capability highlights the practical application of our approach for early detection and intervention in online radicalization. All code and implementation details will be released publicly following the paper’s publication, and data will be made available upon request.
@@ 2 Expanding the Psychometric Toolkit in the Context of Extremism
Over the past 80 years, psychologists and social theorists have extensively explored the psychological and social factors related to extremist beliefs and behaviors. One of the earliest and most influential contributions to the field was Theodor Adorno’s “F-scale,” developed in the aftermath of World War II to measure authoritarian tendencies in individuals  (Adorno et al. 1950). The F-scale sought to better quantify an individual’s susceptibility to fascist and authoritarian attitudes, suggesting that individuals with certain personality traits — such as rigid adherence to conventional values, submission to authority, and aggression toward out-groups — were more likely to adopt extremist ideologies.
@@ 3 Integrating NLP with Social Psychological Theory for Enhanced Extremism Research
Building on this, Hartung et al. (2017) modeled right-wing extremist behaviors by analyzing Twitter data. Their method involved measuring the similarity between user profiles and known extremist or non-extremist groups, offering a nuanced view of how individuals align with extremist ideologies. Unlike traditional binary classification systems, their approach used a continuous scale, allowing for a more dynamic understanding of how individuals engage with extremist content over time. However, the limitation of relying on hand-crafted features and predefined extremist profiles pointed to the need for more flexible models that could capture the evolution of extremism across diverse groups.
Other studies have focused on specific ideological communities, offering more detailed insights into how language fosters radicalization. For example, Ribeiro et al. (2020) analyzed the language of incels (involuntary celibates) and other communities in the “manosphere,” tracing the pathways through which users transition between more moderate and extreme subgroups. This large-scale analysis of online forums revealed how the use of hostile and misogynistic language serves as a gateway to more radical ideologies. Additionally, Dragos et al. (2022) examined the role of emotion in extremist discourse, correlating human annotations of emotion (e.g., anger, hatred) with judgments of whether content was extremist or not. Their findings underscored the importance of emotional language in radicalizing individuals and justifying extremist actions.
@@
Our work highlights opportunities bridge these two approaches — computational analysis and psychological theory — to push our understanding of extremism further. While NLP can effectively identify patterns in language and behavior at scale, wedding these methods with psychological theories of the emotional, cognitive, and social vulnerabilities that drive individuals toward extremism allows for a more holistic approach. This interdisciplinary synthesis enables the development of models that not only detect who is participating in extremist discourse but also explain why individuals are drawn to these movements, offering deeper insights into the underlying mechanisms of radicalization.
@@ 4 Data
We study the unifying characteristics of extremism across various ideological groups, especially those that have been associated with real-world violent attacks. We include (1) posts from the white supremacist/neo-Nazi forum Stormfront; (2) posts from incel communities on Reddit; and (3) quotes from religious or political leaders published in ISIS periodicals (i.e., Dabiq and Rubiyah). To distinguish between the general political discourse and extremist discourse, we include a subset of subreddits from the Politosphere dataset – including multiple subreddits that have since been banned. Further, to contrast the general rhetoric on social platforms and how they compare against the extremist discourse online, we also collected general posts from a set of users on Reddit selected at random. We removed non-English posts, posts with fewer than 10 words, and posts before 2010-01-01 for a more consistent time range since much of the community was not as active and did not exist before then.
@@
2. Incel Reddit (IR) Incels participate in online communities where misogyny and calls for violence against women are prevalent, fueled by the belief that women are withholding a perceived “right” to sex. We collected comments and posts from three subreddits: r/Incels, r/Braincels, and r/Trufemcels, with posts ranging from 2016-07-01 to 2020-04-30. This timeline includes the period when r/Incels was banned (November 2017) for violating Reddit’s policy against incitement of violence toward women, and the subsequent creation of r/Braincels, which gained increasing popularity throughout 2018. The availability and volume of data also allowed us to examine users who began interacting with these forums during this period (§7.3). In total, we collected 51,266 posts from 1,271 users (34,259 posts on r/Braincels, 14,178 posts on r/Incels, and 2,829 posts on r/Trufemcels), with an average post length of $109\pm 160$ words. The IQR was 30-129 words. These three forums were chosen for their high volume of posts and to facilitate a comparison between incels and femcels, who use similar extremist narratives despite being on opposing sides of the ideological spectrum.
3. ISIS Articles Given limited accessibility of English-speaking Islamic extremism discourse on internet forums, we use a small dataset consisting of religious and ideological quotes from two ISIS-supporting magazines: Dabiq and Rumiyah, used to promote their propaganda (Fifth Tribe 2017). It includes scraped data from 15 issues of Dabiq (6/2014 to 7/2016) and 9 issues of Rumiyah (9/2016 to 5/2017), resulting in 2,685 texts. The average message length was $70\pm 127$ and the IQR is 22–71 words. Since the dataset consists solely of quotes from religious and political leaders, no user-level analysis was conducted. Instead, it was used exclusively to learn relevant factors related to Islamic extremism.
@@
2. Politosphere Reddit (PR) We collected comments and posts from the Politosphere Reddit dataset (Hofmann, Schütze, and Pierrehumbert 2022), which consists of subreddits focused on political discourse. This dataset provides us with two unique analysis opportunities. First, many of the items on the extremism scales are inherently or explicitly political in nature, as shown in Table 1. Since extremism is often studied in politico-religious contexts, and both extremist and general political discourse may share similar lexical features, this dataset provides a valuable point of comparison. Second, the meta-data of Politosphere indicates whether the subreddit has been banned, which allows us to explore whether and how extremist factors are present in communities that have violated Reddit policies. Therefore, we sampled subreddits that have been banned as well as those that have not, to explore whether banned subreddits exhibit extremist characteristics. From the collection of 605 subreddits, we randomly selected nine banned and nine not-banned subreddits (see §A.4) resulting in 591,066 posts, with an average post length of $78\pm 113$ words. The IQR is 22-87 words.
@@ 5 Extremism Scales
Our method described in §6 expands the reach and applicability of existing psychological measures designed to quantify traits linked to extremism (Obaidi et al. 2022). Some of these scales were developed to measure specific extremist ideologies, while others focus on more broadly harmful social traits. In this work, we combine all relevant indicators from each scale to create a unified model of extremism factors. In total, we applied 89 items from 10 distinct measurement scales (see Table 1). Our goal was to develop a broadband framework that encompasses a wide range of constructs, spanning past research across various frameworks ranging from general models of extremism to specific ideologies, personality traits, and factors like nationalism and ethnic intolerance. Since these indicators originate from direct self-report questionnaires, all items are phrased as first-person statements, allowing individuals to rate their level of agreement.
@@ 6 Method
Our method is comprised of a Scale Score Extraction Module and a Factor Analysis Module illustrated in Figure  2.
Scale Score Extraction. To score the language content based on the 89 extremism scale items, we employed a method that computes cosine similarity between the vector representations of the post content and each extremism item. This technique, variously referred to as “archetypes” (Varadarajan et al. 2024) or contextualized construct representation (Atari, Omrani, and Dehghani 2023), allows us to quantify the alignment of posts with extremism-related traits. We encode both the posts and survey items using the mxbai-embed-large-v1 model from MixedBread (Lee et al. 2024; Li and Li 2023) using the default hyperparameters. To control for variations in post length, we split each post into 100-word chunks, reducing the impact of length on similarity across our cross-domain datasets (§4).The chunk-level scores are then mean-aggregated to produce post-level extremism item scores.
Exploratory Factor Analysis-based Scoring. Questionnaire scale items are typically factor-analyzed to identify distinct, interpretable factors representing the latent constructs being measured (Shrestha 2021).We apply exploratory factor analysis (EFA) to the 89 extremism scale items pooled together, across the all the datasets combined: White Supremacist Forums, Incel Reddit, ISIS Articles, General Reddit and Politosphere Reddit, resulting in 882,042 data points.
We conduct three key assessments of the suitability of our analysis. First, the Kaiser–Meyer–Olkin (KMO) measure (Kaiser et al. 1974) was used to evaluate sampling adequacy based on the correlation matrix, yielding a KMO value of 0.926, indicating strong suitability for factor analysis. Second, Bartlett’s Test of Sphericity (Gorsuch 1973) was used to assess whether the correlation matrix differs significantly from an identity matrix, with a resulting $p$-value $<$ .001, confirming the appropriateness of the dataset for EFA. Finally, we used Horn’s parallel analysis (Horn 1965) to determine the optimal number of factors for EFA, which was revealed to be 11 factors.
@@ 7.1 Discovered Factors
“Exactly! We do not gain anything by supporting Iran ourselves. It would be much better for everyone if Israel and Iran, devoid of outside assistance or interference, both slaughtered each other
I am strongly against ANY foreign aid (except perhaps White nations), but I believe that giving aid to Israel only is even worse than the debt created by giving aid to all ME countries.”
@@
“They’re moving production of Oreos to Mexico to cut costs anyway. No thanks, I’d rather support American jobs.
@@
“As long as members have veto powers Scotland will never be allowed in.
Table 2: Characteristic differences in posts that signal National Self-Interest in extremist versus general groups.
@@
Starting with RQ1, our goal was to examine how different psychological and social factors contribute to and define radicalization across various extremist groups. We developed an unsupervised model to analyze extremism in online media by comparing the language in posts to established psychological survey scales, followed by factor analysis to score the results. This approach led to the identification of 11 distinct dimensions, which we refer to as “The Extremist Eleven.” Table A1 lists the names of these dimensions, derived from analyzing the content most strongly linked to each factor. These dimensions capture key psychological and social traits commonly associated with extremist ideologies, such as revolutionary attitudes and tendencies toward radical and violent actions.
To assess the face validity of our model, we compared how these factors manifest in extremist communities (White Supremacist, Incel Reddit, and ISIS Articles) versus the broader General Reddit discourse. Figure 3 shows the average factor scores for each group. Positive bars indicate that the factor is a prominent feature of that group, while negative bars indicate its absence. At first glance, the model demonstrates strong face validity. For example, violent groups like White Supremacists (WS) and ISIS score higher on factors such as revolutionary attitudes and radical or violent intentions. We also observe distinctions between WS and ISIS, with the former scoring higher on Social Dominance Orientation and Nationalism, and the latter on State Control and Dogmatism. In contrast, General Reddit shows no significant presence of these extremist factors, except for National Self-Interest. Upon closer examination, this could be explained by the proportionally frequent discussions on news and current affairs in General Reddit, as opposed to the more extreme rhetoric found in the extremist groups. This factor appears in 60% of General Reddit posts, compared to 38% in White Supremacist posts. Table 3 provides examples of the highest-scoring posts for this factor, illustrating how extreme discourse, characteristic of White Supremacist forums, is absent in the top General Reddit posts.
@@
7.2 Forum-level Characterization
We analyze the emergence of the Extremist Eleven factors in communities representing different extremist ideologies in order to identify whether certain psychosocial factors are common across ideologies and whether they reveal characteristics of different forms of extremism (RQ2).
Figure 4: Based on the aforementioned heuristic, we focus only on factor scores $>$ 0. The Politosphere dataset includes a wide range of political forums, often engaging in debates and diverse viewpoints. While many of these forums involve standard political discourse, some show signs of developing extremist tendencies, as observed here. The average factor score across the forums is relatively low, but subreddits like r/debateAltRight score slightly higher. Notably, banned subreddits show significantly elevated scores on dimensions related to radical and violent intent.
We observe different flavors of extremism based on those factors associated with different forums. As observed above (Figure 3), ISIS scores relatively higher in state control and dogmatism, factors that relate to the enforcement of ideology-rooted morals, whereas white supremacist forums have relatively higher signals of social dominance, nationalism, and xenophobia, which could align with Alt-right sentiments. Figure 4, which illustrates the factors in specific the political discourse subreddits in Politosphere, shows these factors as more relevant to Alt-Right communities. Meanwhile, Incel groups exhibit a distinctive psychological profile compared to other extremist groups, marked by higher levels of moral detachment and cold, calculating interpersonal traits. This suggests that, unlike groups driven by ideological zeal or collective identity, incels tend to view social interactions and moral considerations in an emotionally detached, utilitarian way. They are more likely to rationalize harmful attitudes and behaviors without guilt or empathy, framing their grievances in personal rather than overtly political or ideological terms. While other extremist groups may be fueled by a sense of moral superiority or ideological righteousness, incels appear to exhibit a more self-serving and emotionally disengaged approach (see, e.g., Wiggins and Pincus 1989), focusing on personal victimhood and perceived social injustices without regard for broader ethical consequences. This distinct profile positions incels as less ideologically driven but more inwardly focused on their own grievances and frustrations, which they justify through a detached, often hostile worldview.
@@ 7.2 Forum-level Characterization
Next, we want to observe how our model operates on a broader corpus of political discourse text without distinguishing extremist-leaning communities to see how our approach could be used to identify areas with warning signals of extremist discourse. In Figure 4, we analyze the factors in the samples of banned and not-banned subreddits from Politosphere. A key preliminary observation is that there is indeed more signal for radical and violent intent among the banned subreddits. We performed student’s t-tests on the factor distributions between the banned and non-banned subreddits, confirming that the means are significantly different across al the factors. Future work could analyze how the factors change over time, leading up to the point at which they were banned. For example, the r/donaldtrump subreddit was banned after an increase violent rhetoric during the mobilization of the events at the U.S. Capitol on January 6th, 2020 (Isaac and Conger 2021). Analyses similar to the ones described in §7.3, could leverage our model to find changes in extremism signals leading up to this and similar events.
7.3 User-level Analysis
@@
We further analyze two applications of our method to understand the evolution of extremism at a user-level in the context of online communities, where joining or engaging for the first time with an extremist community can be a proxy for radicalization in real life. We compare the Extremist-Eleven Scores of 100 General Reddit users who newly engaged with a forum at some point, and 82 Incel users who were engaged for at least 10 out of 12 consecutive months in the incel community.
@@
Figure A1 shows the performance of a logistic regression model on the average Extremist-Eleven scores derived from posts up until a certain point in time to predict joining the incel forum in the future. To this end, we collected the history of posts made by the 82 incel users in the 12 months leading up to joining the forum. For the general users, we use the time when they post on any new forum (except the incel community) as the time of joining. We perform 5-fold stratified cross validation, reporting the averages of the ROC-AUC across the five runs. Figure A1 demonstrates that individuals who would later go on to engage with the incel community exhibited higher scores on extremist traits well before joining. These elevated scores, up to 12 months prior to joining, reliably differentiate users who are at risk of radicalization long before they actively affiliate with such communities with increasingly high accuracy. This result aligns with theoretical frameworks, such as the  Adorno et al. (1950) work on authoritarianism, which aimed to identify individuals susceptible to extremist ideologies rather than those who were already fully radicalized. Our findings support the notion that personality traits — in this case, for example, moral disengagement and being interpersonally cold and calculating — are reflected in language use, making it possible to predict who might be drawn into these groups well in advance of actual participation.

### v2

@@ Abstract
The proliferation of ideological movements into extremist factions via social media has become a global concern. While radicalization has been studied extensively within the context of specific ideologies, our ability to accurately characterize extremism in more generalizable terms remains underdeveloped. In this paper, we propose a novel method for extracting and analyzing extremist discourse across a range of online ideological community forums. By focusing on verbal behavioral signatures of extremist traits, we develop a framework for quantifying extremism at both user and community levels. Our research identifies 11 distinct factors, which we term “The Extremist Eleven,” as a generalized psychosocial model of extremism. Applying our method to various online communities, we demonstrate an ability to characterize ideologically diverse communities across the 11 extremist traits. We demonstrate the power of this method by analyzing user histories from members of the incel community. We find that our framework accurately predicts which users join the incel community up to 10 months before their actual entry with an AUC of >0.6>0.6, steadily increasing to AUC $\sim 0.9$ three to four months before the event. Further, we find that upon entry into an ideological forum, the users tend to maintain their level of extremist traits within the community, while still remaining distinguishable from the general online discourse. Our findings contribute to the study of extremism by introducing a more holistic, cross-ideological approach that transcends traditional, trait-specific models.
Corresponding Author: Ryan L. Boyd, boyd@utdallas.edu
Codehttps://github.com/humanlab/extremism
@@ 1 Introduction
RQ1. Can an unsupervised method combining NLP techniques and psychosocial theories of extremism identify characteristics of ideological extremism in online communities?
@@
RQ3. Can our framework reliably signal the likelihood of an individual’s eventual active engagement with an extreme ideological community based on their evolving language patterns?
We aim to explore and understand how various psychological and social components contribute to — and characterize — radicalization across a range of extreme ideological groups.  [^1] See §A for semantic clarification. Given that numerous theories have been developed and applied to particular extremist groups and belief systems, we explore a method to unify them in a single model, allowing us to extract factors of extremism in broad, unsupervised data. To make use of existing psychometric and survey-based tools for measuring an individual’s tendency toward extremism at a large scale on social media data, we leverage natural language processing (NLP) techniques to obtain scores for the items using only the available text, requiring no manual labeling or supervision. We identify 11 orthogonal dimensions — the “Extremist Eleven” — that reflect important factors that can be used to broadly characterize key traits linked to extremist ideologies (RQ1). These dimensions are psychosocially meaningful and offer a more complete understanding of the cognitive and emotional drivers that underlie engagement with extremist groups. We explore the face validity of our method by examining how these factors emerge in communities known for extreme-ideological discourse, and in a diverse representation of social media text that captures political discourse more broadly, as well as non-topic-specific discourse.
Additionally, we demonstrate the utility of our methods and results in two preliminary ways. First, we show how the Extremist Eleven can be applied to characterize the psychological composition of both new and existing online communities, offering insights into the shared and unique traits that define these spaces (RQ2). Second, we find that we can accurately and reliably identify individuals who are likely to affiliate with and bind themselves to an extremist online community well in advance — more than 6 months prior to actively joining such communities (RQ3). This predictive capability highlights the practical application of our approach for early detection and intervention in online radicalization. The code and implementation details are available to the public and data can be made available upon request.
@@ 2 Expanding the Psychometric Toolkit in the Context of Extremism
Over the past 80 years, psychologists and social theorists have extensively explored the psychological and social factors related to extremist beliefs and behaviors. One of the earliest and most influential contributions to the field was Theodor Adorno’s “F-scale,” developed in the aftermath of World War II to measure authoritarian tendencies in individuals  (Adorno et al. 1950). The F-scale sought to better quantify an individual’s susceptibility to authoritarian attitudes, suggesting that individuals with certain personality traits — such as rigid adherence to conventional values, submission to authority, and aggression toward out-groups — were more likely to adopt extreme ideologies.
@@ 3 Integrating NLP with Social Psychological Theory for Enhanced Extremism Research
Building on this, Hartung et al. (2017) modeled right-wing extremist behaviors by analyzing Twitter data. Their method involved measuring the similarity between user profiles and known extreme or non-extreme ideological groups, offering a nuanced view of how individuals align with extremist ideologies. Unlike traditional binary classification systems, their approach used a continuous scale, allowing for a more dynamic understanding of how individuals engage with extremist content over time. However, the limitation of relying on hand-crafted features and predefined extremist profiles pointed to the need for more flexible models that could capture the evolution of extremism across diverse groups.
Other studies have focused on specific ideological communities, offering more detailed insights into how language fosters radicalization. For example, Ribeiro et al. (2020) analyzed the language of incels (involuntary celibates) and other communities in the “manosphere,” tracing the pathways through which users transition between more moderate and extreme subgroups. This large-scale analysis of online forums revealed how the use of hostile and misogynistic language serves as a gateway to more radical ideologies. Additionally, Dragos et al. (2022) examined the role of emotion in extremist discourse, correlating human annotations of emotion (e.g., anger, hatred) with judgments of whether the content was extremist or not. Their findings underscored the importance of emotional language in radicalizing individuals and justifying extremist actions.
@@
Our work highlights opportunities to bridge these two approaches — computational analysis and psychological theory — to push our understanding of extremism further. While NLP can effectively identify patterns in language and behavior at scale, wedding these methods with psychological theories of the emotional, cognitive, and social vulnerabilities that drive individuals toward extremism allows for a more holistic approach. This interdisciplinary synthesis enables the development of models that not only detect who is participating in extremist discourse but also explain why individuals are drawn to these movements, offering deeper insights into the underlying mechanisms of radicalization.
@@ 4 Data
We study the unifying characteristics of extremism across various ideological groups, especially those that have been associated with real-world violent attacks. We include (1) posts from the white supremacist/neo-Nazi forum Stormfront; (2) posts from incel communities on Reddit; and (3) quotes from religious or political leaders published in ISIS periodicals (i.e., Dabiq and Rubiyah). To distinguish between general political discourse and extremist discourse, we include a subset of subreddits from the Politosphere dataset – including multiple subreddits that have since been banned. Further, to contrast the general rhetoric on social platforms and how they compare against the extremist discourse online, we also collected general posts from a set of users on Reddit selected at random. We removed non-English posts, posts with fewer than 10 words, and posts before 2010-01-01 for a more consistent time range since much of the community was not as active and did not exist before then.
@@
2. Incel Reddit (IR) Incels participate in online communities where misogyny and calls for violence against women are prevalent, fueled by the belief that women are withholding a perceived “right” to sex. We collected comments and posts from three subreddits: r/Incels, r/Braincels, and r/Trufemcels, with posts ranging from 2016-07-01 to 2020-04-30. This timeline includes the period when r/Incels was banned (November 2017) for violating Reddit’s policy against incitement of violence toward women, and the subsequent creation of r/Braincels, which gained increasing popularity throughout 2018. The availability and volume of data also allowed us to examine users who began interacting with these forums during this period (§7). In total, we collected 51,266 posts from 1,271 users (34,259 posts on r/Braincels, 14,178 posts on r/Incels, and 2,829 posts on r/Trufemcels), with an average post length of $109\pm 160$ words. The IQR was 30-129 words. These three forums were chosen for their high volume of posts and to facilitate a comparison between incels and femcels, who use similar extremist narratives despite being on opposing sides of the ideological spectrum.
3. ISIS Articles Given limited accessibility of English-speaking Islamic extremism discourse on internet forums, we use a dataset consisting of religious and ideological quotes from two ISIS-supporting magazines: Dabiq and Rumiyah, used to promote their propaganda (Fifth Tribe 2017). It includes scraped data from 15 issues of Dabiq (6/2014 to 7/2016) and 9 issues of Rumiyah (9/2016 to 5/2017), resulting in 2,685 texts. The average message length was $70\pm 127$ and the IQR is 22–71 words. Since the dataset consists solely of quotes from religious and political leaders, no user-level analysis was conducted. Instead, it was used exclusively to learn relevant factors related to Islamic extremism.
@@
2. Politosphere Reddit (PR) We collected comments and posts from the Politosphere Reddit dataset (Hofmann, Schütze, and Pierrehumbert 2022), which consists of subreddits focused on political discourse. This dataset provides us with two unique analysis opportunities. First, many of the items on the extremism scales are inherently or explicitly political in nature, as shown in Table 1. Since extremism is often studied in politico-religious contexts, and both extremist and general political discourse may share similar lexical features, this dataset provides a valuable point of comparison. Second, the meta-data of Politosphere indicates whether the subreddit has been banned, which allows us to explore whether and how extremist factors are present in communities that have violated Reddit policies. Therefore, we sampled subreddits that have been banned as well as those that have not, to explore whether banned subreddits exhibit extremist characteristics. From the collection of 605 subreddits, we randomly selected nine banned and nine not-banned subreddits (see §B) resulting in 591,066 posts, with an average post length of $78\pm 113$ words. The IQR is 22-87 words.
@@ 5 Extremism Scales
Our method expands the reach and applicability of existing psychological measures designed to quantify traits linked to extremism (Obaidi et al. 2022). Some of these scales were developed to measure specific extremist ideologies, while others focus on more broadly harmful social traits. In this work, we combine all relevant indicators from each scale to create a unified model of extremism factors. In total, we applied 89 items from 10 distinct measurement scales (see Table 1). Our goal was to develop a broadband framework that encompasses a wide range of constructs, spanning past research across various frameworks ranging from general models of extremism to specific ideologies, personality traits, and factors like nationalism and ethnic intolerance. Since these indicators originate from direct self-report questionnaires, all items are phrased as first-person statements, allowing individuals to rate their level of agreement.
@@ 6 Method
Our method is comprised of a Scale Score Extraction Module and a Factor Analysis Module illustrated in Figure 2.
Scale Score Extraction.
To score the language content based on the 89 extremism scale items, we employed a method that computes cosine similarity between the vector representations of the post content and each extremism item. This technique, variously referred to as “archetypes” (Varadarajan et al. 2024) or contextualized construct representation (Atari, Omrani, and Dehghani 2023), allows us to quantify the alignment of posts with extremism-related traits. We encode both the posts and survey items using the mxbai-embed-large-v1 model from MixedBread (Lee et al. 2024; Li and Li 2023) using the default hyperparameters. To control for variations in post length, we split each post into 100-word chunks, reducing the impact of length on similarity across our cross-domain datasets (§4). The chunk-level scores are then mean-aggregated to produce post-level extremism item scores.
Exploratory Factor Analysis-based Scoring.
Questionnaire scale items are typically factor-analyzed to identify distinct, interpretable factors representing the latent constructs being measured (Shrestha 2021). We apply exploratory factor analysis (EFA) to the 89 extremism scale items pooled together across all the datasets combined: White Supremacist Forums, Incel Reddit, ISIS Articles, General Reddit, and Politosphere Reddit, resulting in 882,042 data points.
We conduct three key assessments of the suitability of our analysis. First, the Kaiser–Meyer–Olkin (KMO) measure (Kaiser et al. 1974) was used to evaluate sampling adequacy based on the correlation matrix, yielding a KMO value of 0.926, indicating strong suitability for factor analysis. Second, Bartlett’s Test of Sphericity (Gorsuch 1973) was used to assess whether the correlation matrix differs significantly from an identity matrix, with a resulting $p$-value << .001, confirming the appropriateness of the dataset for EFA. Finally, we used Horn’s parallel analysis (Horn 1965) to determine the optimal number of factors for EFA, which was revealed to be 11 factors.
@@ 7 Results and Discussion
“Exactly! We do not gain anything by supporting [COUNTRY] ourselves. It would be much better for everyone if [COUNTRY] and [COUNTRY], devoid of outside assistance or interference, both slaughtered each other
I am strongly against ANY foreign aid(except perhaps White nations), but I believe that giving aid to [COUNTRY] only is even worse than the debt created by giving aid to all ME countries.”
@@
“They’re moving production of Oreos to [COUNTRY] to cut costs anyway. No thanks, I’d rather support American jobs.
@@
Table 2: Characteristic differences in posts that signal National Self-Interest in ideological versus general groups.
@@
Starting with RQ1, our goal was to examine how different psychological and social factors contribute to and define radicalization across various extremist groups. We developed an unsupervised model to analyze extremism in online media by comparing the language in posts to established psychological survey scales, followed by factor analysis to score the results. This approach led to the identification of 11 distinct dimensions, which we refer to as “The Extremist Eleven.” Table 4 lists the names of these dimensions, derived from analyzing the content most strongly linked to each factor. These dimensions capture key psychological and social traits commonly associated with extremist ideologies, such as revolutionary attitudes and tendencies toward radical and violent actions.
To assess the face validity of our model, we compared how these factors manifest in extremist communities (White Supremacist, Incel Reddit, and ISIS Articles) versus the broader General Reddit discourse. Figure 3 shows the average factor scores for each group. Positive bars indicate that the factor is a prominent feature of that group, while negative bars indicate its absence. At first glance, the model demonstrates strong face validity. For example, violent groups like White Supremacists (WS) and ISIS score higher on factors such as revolutionary attitudes and radical or violent intentions. We also observe distinctions between WS and ISIS, with the former scoring higher on Social Dominance Orientation and Nationalism, and the latter on State Control and Dogmatism. In contrast, General Reddit shows no significant presence of these extremist factors, except for National Self-Interest. Upon closer examination, this could be explained by the proportionally frequent discussions on news and current affairs in General Reddit, as opposed to the more extreme rhetoric found in extremist groups. This factor appears in 60% of General Reddit posts, compared to 38% in White Supremacist posts. Table 2 provides examples of the highest-scoring posts for this factor, illustrating how extreme discourse, characteristic of White Supremacist forums, is absent in the top General Reddit posts.
@@
Forum-level Characterization
We analyze the emergence of the Extremist Eleven factors in communities representing different extremist ideologies to identify whether certain psychosocial factors are common across ideologies and whether they reveal characteristics of different forms of extremism (RQ2).
Figure 4: Based on the aforementioned heuristic, we focus only on factor scores >> 0. The Politosphere dataset includes a wide range of political forums, often engaging in debates and diverse viewpoints. While many of these forums involve standard political discourse, some show signs of developing extremist tendencies, as observed here. Notably, banned subreddits show significantly elevated scores on dimensions related to radical and violent intent.
We observe different flavors of extremism based on those factors associated with different forums. As observed in Figure 3, ISIS scores relatively higher in state control and dogmatism, factors that relate to the enforcement of ideology-rooted morals, whereas white supremacist forums have relatively higher signals of social dominance, nationalism, and xenophobia, which could align with Alt-right sentiments. Figure 4 illustrates the factors in specific the political discourse subreddits in Politosphere. Meanwhile, Incel groups exhibit a distinctive psychological profile compared to other extremist groups, marked by higher levels of moral detachment and cold, calculating interpersonal traits. This suggests that, unlike groups driven by ideological zeal or collective identity, incels tend to view social interactions and moral considerations in an emotionally detached, utilitarian way. They are more likely to rationalize harmful attitudes and behaviors without guilt or empathy, framing their grievances in personal rather than overtly political or ideological terms. While other extremist groups may be fueled by a sense of moral superiority or ideological righteousness, incels appear to exhibit a more self-serving and emotionally disengaged approach (see, e.g., Wiggins and Pincus 1989), focusing on personal victimhood and perceived social injustices without regard for broader ethical consequences. This distinct profile positions incels as less ideologically driven but more inwardly focused on their own grievances and frustrations, which they justify through a detached, often hostile worldview.
@@
Next, we examine how our model functions within a broader corpus of political discourse without distinguishing extremist-leaning communities, for identifying areas that may indicate warning signals of extremist discourse. In Figure 4, we analyze the factors in samples from both banned and non-banned subreddits within Politosphere. A key preliminary observation is that there is a stronger signal for Radical & Violent Intent and Xenophobia among the banned subreddits. We performed Student’s t-tests on the factor distributions between the banned and non-banned subreddits, confirming that the means are significantly different across all factors. Future work could analyze how the factors change over time, leading up to the point at which they were banned.
User-level Analysis
@@
We further analyze two applications of our method to understand the evolution of extremism at a user level in the context of online communities, where joining or engaging for the first time with an extremist community can be a proxy for radicalization in real life. We compare the Extremist-Eleven Scores of 100 General Reddit users who newly engaged with a forum at some point, and 82 Incel users who were engaged for at least 10 out of 12 consecutive months in the incel community.
@@
Figure 6 shows the performance of a logistic regression model on the average Extremist-Eleven scores derived from posts up until a certain point in time to predict joining the incel forum in the future. To this end, we collected the history of posts made by the 82 incel users in the 12 months leading up to joining the forum. For the general users, we use the time when they post on any new forum (except the incel community) as the time of joining. We perform 5-fold stratified cross-validation, reporting the averages of the ROC-AUC across the five runs. Figure 6 demonstrates that individuals who would later go on to engage with the incel community exhibited higher scores on extremist traits well before joining. These elevated scores, up to 12 months prior to joining, reliably differentiate users who are at risk of radicalization long before they actively affiliate with such communities with increasingly high accuracy. This result aligns with theoretical frameworks, such as the  Adorno et al. (1950) work on authoritarianism, which aimed to identify individuals susceptible to extremist ideologies rather than those who were already fully radicalized. Our findings support the notion that personality traits — in this case, for example, moral disengagement and being interpersonally cold and calculating — are reflected in language use, making it possible to predict who might be drawn into these groups well in advance of actual participation.